OpenStudy (anonymous):

Find the derivative. f(x)=3x^5tan(x)

OpenStudy (across):

\[f(x)=3x^{5\tan(x)}\]or\[f(x)=3x^5\tan(x)\]

OpenStudy (anonymous):

The second one

OpenStudy (across):

\[\frac{df}{dx}=\frac{d}{dx}\left [ 3x^5 \right ]\tan(x)+3x^5\frac{d}{dx}\left [ \tan(x) \right ]\]\[\implies 15x^4\tan(x)+3x^5\sec^2(x).\]This is called the product rule.

OpenStudy (anonymous):

Thamks! That's helpful.